Assume newborn baby weights are Normally distributed with a mean of 7.6lbs and a variance of 0.6lbs. You take a random sample of 3 newborn babies and weight them. Let X be the number of newborn babies in the sample that weight less than 8lbs. Fund Var(X)
Answer #1

Solution :

Given that ,

mean = = 7.6

standard deviation = = 0.61 = 0.7810

n = 3

=   = 7.6

= / n = 0.7810/ 3 = 0.4509

P( < 8) = P(( - ) / < (8-7.6) /0.4509 )

= P(z <0.89 )

= 0.8133

probability = 0.8133